OpenStudy (anonymous):

Jose is paying for a six dollar meal using bills in his wallet. He has four one dollar bills, three five dollar bills, and two ten dollar bills. If he selects two bills at random, one at a time from his wallet, what is the probability that he will choose a one dollar bill and a five dollar bill to pay for the meal? Show your work.

OpenStudy (anonymous):

mathstudent55 please help me

OpenStudy (anonymous):

hello mathstudent55 are you still there

OpenStudy (anonymous):

Jose has 4 ones, 2 fives, and 2 tens. Total of 8 dollar bills. 1st pick = 4 ones / 8 total 2nd pick = 3 ones/7 total answer = 4/8 * 3/7 = about 0.214 = so 21.4%

OpenStudy (anonymous):

Understand @teambringit ?

OpenStudy (anonymous):

he has 3 fives

OpenStudy (anonymous):

never mind thank you anyways

OpenStudy (anonymous):

yw and may I have a medal please :)

OpenStudy (mathstudent55):

4 $1 bills 3 $5 bills 2 $10 bills Total of 9 bills He needs to pick (A) a $1 bill followed by a $5 bill or (B) a $5 followed by a $1 bill. (A) a $1 bill followed by a $5 bill First pick: P($1 bill) = \( \dfrac{4}{ 9} \) Second pick: p($5 bill) = \( \dfrac{3}{8} \) P($1 bill followed by $5 bill) = \(\dfrac{4}{9} × \dfrac{3}{8} = \dfrac{1}{6}\) (B) a $5 bill followed by a $1 bill First pick: P($5 bill) = \( \dfrac{3}{9} \) Second pick: P($1 bill) = \( \dfrac{4}{8} \) P($5 bill followed by $1 bill) = \(\dfrac{3}{9} × \dfrac{4}{8} = \dfrac{1}{6}\) P($5 and $1 in either order) = \(\dfrac{1}{6} + \dfrac{1}{6} = \dfrac{2}{6} = \dfrac{1}{3} \)